OpenStudy (anonymous):

7x - 3/6 = 6x - 5/8?

OpenStudy (anonymous):

Put all x in one side and all the constants in the other side. \[7x - \frac36 =6x - \frac58\] \[7x-(6x)-\frac36=-\frac58\] \[7x-(6x)=-\frac58+\frac36\] x = -1/8